# FuSeFL: Fully Secure and Scalable Cross-Silo Federated Learning

### v2

@@ IV-A Notations and Preliminaries
$D_{i}]_{j}$ $j^{th}$ share of the secret shared data of $i^{th}$ client
@@ IV-B Federated Learning (FL)
Traditional machine learning requires aggregating data from multiple sources onto a central server for training. However, this approach is often impractical due to privacy regulations, communication costs, and bandwidth constraints [2].
Federated Learning (FL) addresses these challenges by enabling collaborative model training across distributed clients without exposing raw data [1, 2]. As illustrated in Fig. 3, (1) the central server first distributes the current global model to a subset of participating clients. (2) Each client trains the model locally using its private dataset. (3) The clients then send their model updates back to the server. (4) The server aggregates these updates (e.g., via FedAvg [2]) to refine the global model. This process is repeated for multiple rounds until convergence.
Figure 3: Overview of Federated Learning. (1) The server sends the global model to the clients. (2) The clients then train the model locally and (3) send back their trained models to the server. (4) The server updates the global model for the next round of training.
FedAvg [2] is the most widely adopted aggregation method in federated learning. It computes the global model by taking a weighted average of the local model parameters ($w$) from participating clients:
@@
Here, $w_{t}^{k}$ denotes the model parameters from client $k$ at round $t$, $n_{k}$ is the number of training samples on client $k$, and $n=\sum_{k=1}^{K}n_{k}$ is the total number of samples across all clients. This weighted averaging ensures that clients with larger datasets have proportionally greater influence on the updated global model.
Cross-silo federated learning targets deployments where each client represents a large, resource-rich organization, such as a hospital or financial institution, with stable availability and ample compute, storage, and bandwidth [8]. While these settings offer stronger infrastructure support than cross-device FL, they also present distinct challenges. The large datasets held by each silo amplify communication and memory demands, resulting in scalability bottlenecks, particularly at the aggregator, where per-client updates must be processed and stored [37, 1, 8].
IV-C Secure Multi-Party Computation (MPC)
MPC is a cryptographic technique that enables multiple parties to collaboratively compute a function over their private inputs while ensuring that the inputs remain confidential and individual shares are never disclosed during the process. By splitting sensitive data into encrypted or secret shares and processing them separately, secure computation is achieved.
MPC allows computation over encrypted data while remaining lightweight, making it a practical option for distributing computation securely across multiple clients without incurring heavy computational costs. However, as the number of participants increases, the communication overhead in MPC protocols also grows significantly.
Additive Sharing. To perform linear operations, additive sharing [38, 39, 40, 41] is a common MPC technique that splits private data into multiple secret shares, each distributed to a different party. In this approach, the data is secret-shared such that no single party can reconstruct it independently. We define additive secret sharing over a finite ring $\mathbb{Z}_{2^{n}}$, where $n$ is the bit-length of encoded values and arithmetic is performed modulo $2^{n}$.
Let $X\in\mathbb{Z}_{2^{n}}$ be a private value. In a setting with j parties, $X$ is split into j additive shares ($x]_{j}$) and is represented as the sum of its shares:
@@ IV-C Secure Multi-Party Computation (MPC)
  • Multiplication: To perform multiplication over additive shares, Beaver triples [42] are used. If $Z]=[X]\cdot[Y]$, and the precomputed triple is $c]=[a]\cdot[b] [$22], each party $P_{j}$ does the following:
@@
Function Secret Sharing (FSS). Unlike additive sharing, Function Secret Sharing [43, 44, 45] splits the function $f:\mathbb{Z}_{2^{n}}\rightarrow\mathbb{Z}_{2^{n}}$ across parties. Each party holds a share of the function rather than the data, and computation is done by applying these function shares to the input.
If $f(x)$ is the function to evaluate and $x$ is the (possibly masked) input, then:
@@
Fig. 4(b) illustrates that the private input $X$ is masked as $X_{c}=X+R$, and each party evaluates its function share on $X_{c}$. The final result is reconstructed by summing the outputs.
@@
IV-D AriaNN: Federated Learning Extension
Ryffel et al. [25] proposed AriaNN, which uses additive sharing and function secret sharing (FSS) [43, 45] to privately perform deep learning tasks (both training and inference) with minimal communication overhead. In their scheme, both the data and the model are secret-shared between two parties to ensure that no single party can reconstruct the sensitive information. AriaNN supports several operations required for neural network training and inference.
AriaNN uses additive secret sharing to securely perform linear computations (explained in §IV-C). For example, to support secure multiplication and convolution, AriaNN employs Beaver triples [42], specifically matrix-based Beaver triples [22], for efficient batch operations.
@@ IV-D AriaNN: Federated Learning Extension
To perform comparison-based computations like ReLU, AriaNN uses FSS [43, 45] (explained in §IV-C). By using FSS, AriaNN achieves a significant reduction in the number of communication rounds compared to previous works [40, 23, 24]. To implement comparison $(a\leq\alpha)$, AriaNN uses private equality testing proposed by [45]. The evaluators evaluate all possible paths before the special path (where the public value $a$ is equal to the private value $\alpha$). Then, they sum them up to obtain the final output of the comparison [25].
AriaNN, then, extended their secure training framework to support federated learning, enabling privacy-preserving model training across several distributed clients.
@@
The training process begins by secret-sharing the model parameters between the server and clients, each holds one share of the model. As illustrated in Fig. 5, the protocol proceeds as follows: (1) Each client secret-shares a portion of its local data with the server while retaining the other share locally. (2) Clients and the server then collaboratively train their respective shares of the local model using MPC over a number of training iterations. (3) After local training, one designated client (e.g., $C_{A}$) acts as the aggregator, collecting secret-shared updates from other clients. (4) The aggregator performs secure weighted averaging of these updates, ensuring that individual data and model privacy are preserved throughout. The server, in parallel, aggregates its corresponding shares of the model updates. (5) The updated model share $GM]_{1}$ is redistributed privately by the aggregator to all clients for the next round, while the server also updates its copy $GM]_{0}$ accordingly.
This process repeats for a fixed number of global epochs. At the end of training, the final model is reconstructed by combining the two secret shares:
$GM=[GM]_{0}+[GM]_{1}$
@@
Since the model remains secret-shared throughout training, this design protects both data and model confidentiality while enabling secure collaborative learning.
@@
Figure 5: Overview of the AriaNN-FL scheme. (2) Each client collaboratively trains its local model with the server. The server holds a share of the model and the data, while the remaining shares are securely stored on the client side. (4) $C_{A}$, i.e., client aggregator, aggregates the client-side shares of the local models.
@@ V Fully Secure Federated Learning (FuSeFL)
In this section, we present FuSeFL, a federated learning framework that holistically protects both data and model confidentiality while enabling scalable training in cross-silo environments. FuSeFL addresses key limitations of prior work, including privacy vulnerabilities and centralized bottlenecks. In many existing systems, the server performs per-client training or coordination, limiting parallelism and creating a performance bottleneck. This becomes especially restrictive in cross-silo settings, where clients are powerful and data-rich. FuSeFL overcomes this by decentralizing training across client groups and assigning the server only a lightweight aggregation role, enabling secure and efficient scalability.
@@ V-A Overview of FuSeFL
To enable secure distributed training, FuSeFL leverages multi-party computation to keep both data and model parameters secret-shared throughout. The protocol builds on AriaNN’s MPC backend but rearchitects the system to support federated deployment, enhance scalability, and lower communication latency.
In each training round, clients are grouped into pairs. Within each group, both the model and client data ($D_{k}$ and $D_{L}$) are secret-shared to ensure that no single party can access them in plaintext. As shown in Fig. 6, the workflow proceeds as follows:
@@
    Secure Group Training: Clients $C_{k}$ and $C_{l}$ collaboratively train a local model over their secret-shared data ($D_{k}$ and $D_{L}$) using MPC. In our case, we use AriaNN to perform secure training within a group (§IV-D). The model itself and the data are maintained in a secret-shared form throughout training, ensuring confidentiality of both inputs and intermediate states. One share of the data remains local on each client to ensure that the data remains confidential and never leaves the client in its entirety.
@@
[…]
@@
By structuring the training process around small, secure client groups and decoupling server-side training from aggregation, FuSeFL achieves a practical and scalable solution for secure FL in regulated cross-silo deployments.
@@ V-B Risk-Aware Dynamic Client Grouping
To reduce the likelihood of information leakage through collusion, FuSeFL employs a dynamic, adversary-aware grouping strategy in which client pairs are reshuffled every global round. This continual reconfiguration enhances robustness by minimizing repeated co-training relationships, thereby amplifying privacy through exposure diversity.
Client grouping is coordinated by the server using only public or registered metadata, such as organizational type, legal jurisdiction, or historical affiliations, while avoiding access to private data or training artifacts. This design choice supports deployment in cross-silo settings where participants (e.g., hospitals, banks, or labs) operate under distinct governance and regulatory constraints.
Each client pair is assigned a collusion risk score $\rho(C_{i},C_{j})\in[0,1]$, reflecting potential structural correlations such as shared ownership, industry overlap, or regulatory proximity. Given these scores, FuSeFL computes a set of disjoint client pairs $\mathcal{G}=\{(C_{i},C_{j})\}$ that minimize overall collusion exposure:
@@
Establishing MPC Channels: Once groups are selected, each pair $(C_{i},C_{j})$ initiates a secure MPC session using authenticated metadata exchange and ephemeral key agreement. These sessions are short-lived and introduce negligible delay (typically under 30ms per pair [25]). Channels are secured via TLS or mutual-authenticated Diffie-Hellman, in alignment with enterprise-grade deployments.
This risk-aware, round-wise reshuffling reduces long-term collusion opportunities, ensures regulatory separation, and strengthens overall privacy guarantees in multi-institution FL settings.
@@ V-C Training Modes: Serial and Parallel
In FuSeFL-Serial (Fig. 7(a)), each client group jointly trains a single model instance using the data from both clients, one after the other. First, the local model $LM_{1}$ is trained using client $C_{1}$’s data $D_{1}$ under MPC. Then, the same model is further trained using client $C_{2}$’s data $D_{2}$. This sequential training leverages the full dataset available to the group, typically leading to improved model accuracy. Since only one model instance is maintained, this mode is memory-efficient. However, it incurs higher training latency due to its serialized nature. Algorithm 1 provides implementation details.
In contrast, FuSeFL-Parallel (Fig. 7(b)) accelerates training by allowing both clients to train separate local models concurrently. The group maintains two model instances ($LM_{1}$ and $LM_{2}$), which are secret-shared among the clients. Each model is trained independently on $D_{1}$ and $D_{2}$ in parallel. After training, the clients locally aggregate the two updated models before securely transmitting the result to the servers. This mode reduces training latency and is better suited for time-sensitive deployments (evaluated in §VII-B), at the cost of slightly higher memory usage and a modest trade-off in final model accuracy (evaluated in §VII-A). Details are outlined in Algorithm 2).
Both training modes fully preserve data and model confidentiality through MPC, ensuring that no single client or server can reconstruct sensitive information. Table II summarizes the trade-offs between the two modes.
TABLE II: Comparison of FuSeFL-Serial and FuSeFL-Parallel in terms of performance and resource trade-offs.
Aspect FuSeFL-Serial FuSeFL-Parallel
Training Latency Higher Lower
Model Accuracy Slightly higher Slightly lower
Memory Usage Lower Higher
Best Use Case Accuracy- or memory- sensitive applications Latency-sensitive applications
Input: $j\leftarrow 2$
     Global model $GM]_{j}$;
     Local models $LM_{(i/2)}]_{j}$;
     Local datasets $D_{i}]_{j}$
for s in j do
  Server s $\leftarrow$ $GM]_{j}$
end for
for each global epoch do
 for each group g (in parallel) do
    for each client c in g do
        for s in j do
           Server s sends $GM]_{j}$ to client c
        end for
       Client c $\leftarrow$ $D_{c}]_{j}$
        Client c $\leftarrow$ $D_{(j-c)}]_{(j-c)}$
        $LM_{g}]_{j} \leftarrow [$$GM]_{j}$
     end for
    for each local epoch do
        clients train $LM_{g}]$ using $D_{0}]$
     end for
    for each local epoch do
        clients train $LM_{g}]$ using $D_{1}]$
     end for
    for each client c in g do
        for s in j do
           client c sends $LM_{g}]_{j}$ to server s
        end for
     end for
  end for
 $GM]_{j}$ = $AVG(\Sigma{[LM_{g}]_{j})})$
end for
Algorithm 1 FuSeFL-Serial
@@ V-D Communication Complexity
Secure federated learning systems incur communication overhead primarily in two channels: server-to/from-client (S2C/C2S) and client-to-client (C2C).
Server-to-Client / Client-to-Server (S2C/C2S). In AriaNN-FL, the server actively participates in secure training by running MPC protocols jointly with each client across every local epoch. This introduces a severe server-side bottleneck as the number of clients grows. In contrast, FuSeFL decouples training from the server and restricts its role to model aggregation. Clients only communicate their secret-shared updates to two non-colluding servers once per global round. The overall S2C/C2S complexity is reduced to $\mathcal{O}((n/2)\cdot x)$, assuming a model size of $x$ and $n$ total clients. This significantly improves scalability and enables parallel execution. Client-to-Client (C2C). AriaNN-FL performs client-to-client communication during aggregation, where each client sends its model share to a designated aggregator for reconstruction, which incurs $\mathcal{O}(n\cdot x)$ communication and requires one trusted aggregator per round. In contrast, FuSeFL confines secure training to fixed client pairs using MPC. Each pair exchanges secret shares of data and model parameters locally, with communication parallelized across groups. This keeps per-group C2C overhead constant and independent of the total number of clients.
By restructuring the communication pattern to exploit pairwise training and two-server aggregation, FuSeFL substantially reduces both the volume and centralization of communication (evaluated in §VII-C). This architectural change eliminates per-epoch server interaction, enables system-wide parallelism, and improves bandwidth efficiency without weakening security guarantees.

### v3

@@ IV-A Notations and Preliminaries
$GM]_{j}$, $LM_{i}]_{j}$, $D_{i}]_{j}$ $j^{th}$ share of the secret shared data/model of $i^{th}$ client
@@ IV-B Federated Learning (FL)
Traditional machine learning requires aggregating data from multiple sources onto a central server for training. However, this approach is often impractical due to privacy regulations, communication costs, and bandwidth constraints [2]. FL addresses these challenges by enabling collaborative model training across distributed clients without exposing raw data [1, 2]. As illustrated in Fig. 3, (1) the central server first distributes the current global model to a subset of participating clients. (2) Each client trains the model locally using its private dataset. (3) The clients then send their model updates back to the server. (4) The server aggregates these updates (e.g., via FedAvg [2]) to refine the global model. This process is repeated for multiple rounds until convergence.
Figure 3: FL overview. (1) The server sends the global model to the clients. (2) The clients then train the model locally and (3) send back their trained models to the server. (4) The server updates the global model for the next round of training.
FedAvg [2] is the most widely adopted aggregation method in FL. It computes the global model by taking a weighted average of the local model parameters ($w$) from participating clients:
@@
Here, $w_{t}^{k}$ denotes the model parameters from client $k$ at round $t$, $n_{k}$ is the number of training samples on client $k$, and $n=\sum_{k=1}^{K}n_{k}$ is the total number of samples across all clients. This weighted averaging ensures that clients with larger datasets have proportionally greater influence on the updated global model. In FL, each client’s dataset and the model size can amplify communication and memory demands, leading to scalability bottlenecks, particularly at the aggregator, where per-client updates must be processed and stored [37, 1, 8]. This occurs because we must keep a copy of all submitted updates to determine the weight associated with each client before performing the aggregation.
IV-C Mix Networks (mixnets)
Mix networks (mixnets) enable privacy-preserving communication by routing messages through a sequence of intermediate servers that shuffle, re-encrypt, and delay traffic to obscure the link between senders and receivers. This process provides strong anonymity by disrupting traffic analysis and timing correlations, effectively concealing communication patterns from adversaries [38, 39]. Although mixnets introduce additional latency due to multi-hop routing and layered encryption, modern low-latency designs (e.g., Loopix [40], Stadium [41]) significantly mitigate this overhead, achieving delays of only a few seconds while preserving robust anonymity. Consequently, mixnets remain practical for privacy-sensitive applications such as anonymous email systems [42, 43] and electronic voting protocols [44, 45].
IV-D Secure Multi-Party Computation (MPC)
MPC is a cryptographic technique that enables multiple parties to collaboratively compute a function over their private inputs while ensuring that the inputs remain confidential and individual shares are never disclosed during the process. By splitting sensitive data into encrypted or secret shares and processing them separately, secure computation is achieved. MPC allows computation over encrypted data while remaining lightweight, making it a practical option for distributing computation securely across multiple clients without incurring heavy computational costs. However, as the number of participants increases, the communication overhead in MPC protocols also grows significantly.
Additive Sharing. To perform linear operations, additive sharing [46, 47, 48, 49] is a common MPC technique that splits private data into multiple secret shares, each distributed to a different party. In this approach, the data is secret-shared such that no single party can reconstruct it independently. We define additive secret sharing over a finite ring $\mathbb{Z}_{2^{n}}$, where $n$ is the bit-length of encoded values and arithmetic is performed modulo $2^{n}$. Let $X\in\mathbb{Z}_{2^{n}}$ be a private value. In a setting with j parties, $X$ is split into j additive shares ($x]_{j}$) and is represented as the sum of its shares:
@@ IV-D Secure Multi-Party Computation (MPC)
  • Multiplication: To perform multiplication over additive shares, Beaver triples [50] are used. If $Z]=[X]\cdot[Y]$, and the precomputed triple is $c]=[a]\cdot[b] [$19], each party $P_{j}$ does the following:
@@
Function Secret Sharing (FSS). Unlike additive sharing, Function Secret Sharing [51, 52, 53] splits the function $f:\mathbb{Z}_{2^{n}}\rightarrow\mathbb{Z}_{2^{n}}$ across parties. Each party holds a share of the function rather than the data, and computation is done by applying these function shares to the input. If $f(x)$ is the function to evaluate and $x$ is the (possibly masked) input, then:
@@
Fig. 4(b) illustrates that the private input $X$ is masked as $X_{c}=X+R$, and each party evaluates its function share on $X_{c}$. The result is reconstructed by summing the outputs.
@@
IV-E AriaNN: Federated Learning Extension
@@
Ryffel et al. [22] proposed AriaNN, which uses additive sharing and function secret sharing (FSS) [51, 53] to privately perform deep learning tasks (both training and inference) with minimal communication overhead. In their scheme, both the data and the model are secret-shared between two parties to ensure that no single party can reconstruct the sensitive information. AriaNN supports several operations required for neural network training and inference. AriaNN uses additive secret sharing to securely perform linear computations (explained in §IV-D). For example, to support secure multiplication and convolution, AriaNN employs Beaver triples [50], specifically matrix-based Beaver triples [19], for efficient batch operations. To perform comparison-based computations like ReLU, AriaNN uses FSS [51, 53] (explained in §IV-D). By using FSS, AriaNN achieves a significant reduction in the number of communication rounds compared to previous works [48, 20, 21]. To implement comparison $(a\leq\alpha)$, AriaNN uses private equality testing proposed by [53]. The evaluators evaluate all possible paths before the special path (where the public value $a$ is equal to the private value $\alpha$). Then, they sum them up to obtain the final output of the comparison [22].
@@
AriaNN, then, extended their secure training framework to support federated learning, enabling privacy-preserving model training across several distributed clients. The training process begins by secret-sharing the model parameters between the server and clients, each holds one share of the model. As illustrated in Fig. 5, the protocol proceeds as follows: (1) Each client secret-shares a portion of its local data with the server while retaining the other share locally. (2) Clients and the server then collaboratively train their respective shares of the local model using MPC for some training iterations. (3) After local training, one designated client (e.g., $C_{A}$) acts as the aggregator, collecting secret-shared updates from other clients. (4) The aggregator performs secure weighted averaging of these updates, ensuring that individual data and model privacy are preserved throughout. The server, in parallel, aggregates its corresponding shares of the updates. (5) The updated model share $GM]_{1}$ is redistributed privately by the aggregator to all clients for the next round, while the server updates its copy $GM]_{0}$ accordingly.
@@
This process repeats for a fixed number of global epochs. At the end of training, the final model is reconstructed by combining the two secret shares: $GM=[GM]_{0}+[GM]_{1}$ Since the model remains secret-shared throughout training, this design protects both data and model confidentiality while enabling collaborative learning.
@@
Figure 5: AriaNN-FL overview. (2) Each client collaboratively trains its local model with the server. The server holds a share of the model and the data, while the remaining shares are securely stored on the client side. (4) $C_{A}$, i.e., client aggregator, aggregates the client-side shares of the local models.
@@ V Fully Secure Federated Learning (FuSeFL)
In this section, we present FuSeFL, an FL framework that holistically protects both data and model confidentiality while enabling scalable training. FuSeFL addresses key limitations of prior work, including privacy vulnerabilities and centralized bottlenecks. In existing systems, the server performs per-client training or coordination, limiting parallelism and creating a performance bottleneck. This becomes especially restrictive in settings where clients are data-rich. FuSeFL overcomes this by decentralizing training across client groups and assigning the server only lightweight computations for aggregation and group pairing, enabling secure and efficient scalability.
@@ V-A Overview of FuSeFL
To enable secure distributed training, FuSeFL leverages MPC to keep both data and model parameters secret-shared throughout. The protocol builds on AriaNN’s MPC backend but rearchitects the system to support federated deployment, enhance scalability, and lower training time. To prevent collusion among computing parties, we assume that one server is trusted. This trusted server acts as a mix, coordinating the computing clients. All training-related data communication among groups is routed through it, ensuring that both senders and receivers remain anonymous. In each training round, clients are anonymously grouped into pairs. Within each group, both the model and client data ($D_{k}$ and $D_{L}$) are secret-shared to ensure that no single party can access them in plaintext. As shown in Fig. 6, the workflow proceeds as follows:
@@
    Secure Group Training: Clients $C_{k}$ and $C_{l}$ collaboratively train a local model over their secret-shared data ($D_{k}$ and $D_{L}$) using MPC. In our case, we use AriaNN to perform secure training within a group (§IV-E). The model itself and the data are maintained in a secret-shared form throughout training, ensuring confidentiality of both inputs and intermediate states. One share of the data remains local on each client to ensure that the data remains confidential and never leaves the client in its entirety. All data communication between pairs of clients in this stage is routed anonymously through the mix, $S_{1}$, which is the trusted server. $S_{1}$ pairs the clients using the Risk-Aware Dynamic Client Grouping algorithm and routes the training information accordingly. This additional communication layer is used to route, shuffle, and delay the packets in order to keep the pairs anonymous.
@@
    Update Submission: After completing local training for several epochs, each client returns their share of the updated local model to the corresponding server. $C_{k}$ sends $LM]_{1}$ to $S_{1}$ and $C_{l}$ sends $LM]_{2}$ to $S_{2}$.
@@
By structuring the training process around small, secure client groups and decoupling server-side training from aggregation, FuSeFL achieves a practical and scalable solution for secure FL.
@@ V-B Risk-Aware Dynamic Client Grouping
To reduce the likelihood of weakening anonymity through flow-matching attacks[54] and backdoor channels that may lead to information leakage and collusion over time, FuSeFL employs a dynamic, adversary-aware grouping strategy in which client pairs are reshuffled every global round. This continual reconfiguration enhances robustness by minimizing repeated co-training relationships, thereby amplifying privacy through exposure diversity and repeated communication patterns. Each client pair is assigned a collusion risk score $\rho(C_{i},C_{j})\in[0,1]$, reflecting potential anonymity weakening and structural correlations such as shared ownership, industry overlap, or regulatory proximity. Given these scores, FuSeFL computes a set of disjoint client pairs $\mathcal{G}=\{(C_{i},C_{j})\}$ that minimize overall collusion exposure:
@@
Establishing MPC Channels: Once groups are selected, each pair $(C_{i},C_{j})$ anonymously initiates a secure MPC session using authenticated metadata exchange and ephemeral key agreement through the trusted server node. These sessions are short-lived and introduce negligible delay (typically under 30ms per pair [22]). Channels are secured via TLS or mutual-authenticated Diffie-Hellman, in alignment with enterprise-grade deployments.
This risk-aware, round-wise reshuffling reduces long-term collusion opportunities, ensures regulatory separation, and strengthens overall privacy guarantees in FL settings.
@@ V-C Training Modes: Serial and Parallel
In FuSeFL-Serial (Fig. 7(a)), each client group jointly trains a single model instance using the data from both clients, one after the other. First, the local model $LM_{1}$ is trained using client $C_{1}$’s data $D_{1}$ under MPC. Then, the same model is further trained using client $C_{2}$’s data $D_{2}$. This sequential training leverages the full dataset available to the group, leading to improved model accuracy. Since only one model instance is maintained, this mode is memory-efficient. However, it incurs higher training latency due to its serialized nature.
In contrast, FuSeFL-Parallel (Fig. 7(b)) accelerates training by allowing both clients to train separate local models concurrently. The group maintains two model instances ($LM_{1}$ and $LM_{2}$), which are secret-shared among the clients. Each model is trained independently on $D_{1}$ and $D_{2}$ in parallel. After training, the clients locally aggregate the two updated models before securely transmitting the result to the servers. This mode reduces training latency and is better suited for time-sensitive deployments (evaluated in §VII-A), at the cost of slightly higher memory usage and a modest trade-off in final model accuracy (evaluated in §VII-D). Details are outlined in Algorithm 1. Both training modes fully preserve data and model confidentiality through MPC, ensuring that no single client or server can reconstruct sensitive information.
@@ V-D Communication Complexity
Secure FL systems incur communication overhead in two channels: server-to/from-client (S2C/C2S) and client-to-client (C2C).
Server-to-Client / Client-to-Server (S2C/C2S). In AriaNN-FL, the server actively participates in secure training by running MPC protocols jointly with each client across every local epoch. This introduces a severe server-side computational bottleneck as the number of clients grows. In contrast, FuSeFL decouples training from the server and restricts its role to model aggregation, client pairing, and data routing. Clients in each group exchange data through a trusted server, and at the end of each round, they send their secret-shared updates to two non-colluding servers. The overall direct S2C/C2S complexity for aggregation and distribution is reduced to $\mathcal{O}(n\cdot x)$, assuming a model size of $x$ and $n$ total clients. This significantly improves scalability and enables parallel execution. The data communication required to distribute clients’ data, however, increases from $\mathcal{O}(n\cdot d)$ to $\mathcal{O}(n\cdot(2\cdot d))$, assuming that the redirection delay at the server is negligible, where $d$ is the data size of each client and $n$ is the total number of clients.
Client-to-Client (C2C). AriaNN-FL performs client-to-client communication during aggregation and distribution, where each client sends its model share to a designated aggregator for reconstruction and gets back the updated global model, which incurs $\mathcal{O}(2\cdot n\cdot x)$ communication overhead, assuming a model size of $x$ and $n$ total clients. In contrast, FuSeFL confines secure training to fixed client pairs using MPC. Each pair exchanges secret shares of data and model parameters through the trusted server. This increases the per-group C2C overhead to achieve a higher level of security, which is the non-colluding clients. constant and independent of the total number of clients. By restructuring the communication pattern to exploit pairwise training and two-server aggregation, FuSeFL substantially reduces the training time while increasing the communication overhead (evaluated in §VII-A §and VII-B). This architectural change eliminates server bottlenecks and enables system-wide parallelism without weakening security guarantees.